C4H6O6 * 3 H2O (trihydride) is used to prepare V= 400 mL with a concentration of 0.2 mol/L of C4H6O6. How much of pure anhydride C4H6O6 (tartaric acid), is required to make V= 400 mL with a concentration of 0.2 mol/L?

Please show all steps.
Answer #1

Volume (V ) = 400 mL = 0.4 L.

Concentration = 0.2 mol/L.

Then, moles of tartaric acid = volume×concentration

= 0.4×0.2 = 0.08 .

Mass of anhydride tartaric acid

= moles × molar mass = 0.08×150 = 12 g.

Hence, 12 g pure anhydride tartaric acid is required.

[ Note; molar mass of pure anhydride tartaric acid =150 g/mol].